OpenStudy (anonymous):

Quick identity question. Does cot (x/2) = tan (2/x) ?

jimthompson5910 (jim_thompson5910):

no you must be thinking of cot(x) = 1/tan(x) or tan(x) = 1/cot(x)

OpenStudy (anonymous):

what is the density of water in newton per meter cube?

OpenStudy (anonymous):

JohnE95, that question belongs in a different section It also could probably be solved by a google search

hartnn (hartnn):

or \(\cot^{-1}(x/2)=\tan^{-1}(2/x)\) ?